# Inflationary Dynamics in Einstein-Gauss-Bonnet Gravity Using New Slow-Roll Approximations considering Generalised Reheating

### v1

@@ Abstract
The Cosmic Microwave Background (CMB) observations results reported by Planck satellite suggest a plateau characteristic in the flat potential of the inflaton field is favored to drive the Universe’s early acceleration. On that note, Power Law Plateau (PLP) potential has gained a lot of attention. In this study, we demonstrate that implementing this inflationary model in an Einstein-Gauss-Bonnet(EGB) gravity background makes the model compatible with observations while avoiding late-stage thermal inflation which is otherwise unavoidable in a standard scenario. More importantly, we have taken the PLP model as a case study to check the new slow-roll approximations method proposed in [1]. Then we take one step further to check the consequence of generalized reheating in this scenario. Thus opening a new promising window to study inflationary dynamics in EGB gravity.
@@ I Introduction
The inflationary epoch, along with the standard hot Big Bang model, could resolute the understanding of the birth and evolution of our Universe to its current state [2], [3]). The first idea of inflation was proposed by Guth [4] and since then the inflationary epoch has gained tremendous attention and it has become an interdisciplinary study requiring both the fields of high energy physics and cosmology[5, 6, 7, 8, 9, 10]. CMB observations by the Planck and Wilkinson Microwave Anisotropy Probe(WMAP) missions have ruled out or strongly disfavoured most of the textbook models of inflation in the standard scenario[11, 12, 13]. Based on the current observation, it becomes evident that a plateau-like feature in the potential can satisfy the constraints imposed by CMB in the best possible manner. In a key study, Dimopoulos and Owen [14] presented the Power Law Plateau inflation (PLP) with plateau properties. However, in this theoretically justified model, one needs to incorporate a phase of thermal inflation to let it withstand the observational constraints while remaining in the conventional cold inflationary premise.
@@
It was shown for the first time in [15] that, if one wants to develop the inflationary models under the UV complete string theory, PLP in the alternative theory of Randall-Sundrum brane-world[16, 17]can be compatible with the observation, along with it can circumvent the tighter constraints from string theory like swampland conjecture[18]. But, on the other hand, RS models suffer serious existential crises because of the continuous limitation on the scale provided by the braneworld theory[19]. It has been shown that in modified gravity the restriction on the inflationary models can be less stringent[20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34]. In other words, models which are disfavoured by observation in the standard cold inflationary can satisfy the observational constraints in the modified gravity. The addition of the Gauss-Bonnet component to the Einstein-Hilbert action does not affect the equations of motion as it is a total derivative. When EGB coupling $\xi(\phi)$ is combined with a function of the scalar field $\phi$, it becomes dynamically important. In string theory, the Gauss-Bonnet term serves as a quantum correction to the Einstein-Hilbert action. Numerous inflationary models have been studied in this context[35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 52, 67, 68, 69, 70, 71, 72], most widely studied EGB coupling is the inverse function of the scalar field[36, 37, 38, 39, 41, 49, 51, 54, 55]. However in this manuscript we will study different form of coupling which involves a ``$\tanh"$ term[73]. In [1] the authors have presented a new kind of slow-roll approximations, which are found to be more accurate in comparison to the previous slow-roll approximations presented in[45]. As it has been already shown in [1] that new slow approximations are more accurate, we aim to do a case study of the new slow-roll approximations in the EGB scenario using the the PLP potential as a prototype. This serves the primary objective of this exercise. More importantly, we extend the study to accommodate the generalized reheating and review its effect in the effective EGB background. The standard cold inflationary scenario requires an epoch of reheating after the end of inflation to enter the radiation-dominated epoch and start the Big Bang Nucleosynthesis (BBN). Studying the reheating phase following the end of inflation is necessary for the conceptual development of inflationary model construction. We have studied reheating in this work and established the limits on the number of e-folding and reheating temperatures that are permitted by CMB data.
@@
The paper is organized as follows: In section II, we will go over some of the fundamental critical features in the EGB framework. A theoretically well-motivated Power Law Plateau (PLP) potential for this study has been reviewed in section III. In section IV, the calculation of slow-roll parameters without any approximations, is reviewed. In section V, the new slow-roll approximations (SRA) as reported in the work [1] are reviewed. Finally, the analysis of the inflationary dynamics using the new SRA and the effect of generalized reheating has been reported in the sections VI and VII. Finally, we conclude in section VIII with our conclusions and future direction of work.
@@ II Review of Einstein-Gauss Bonnet Gravity
Evolution of equations and dynamical system in Einstein-Gauss-Bonnet gravity can be written for a spatially flat FLRW metric [35, 45]as:
@@
where $H=\dot{a}/a$ denotes the Hubble parameter and $a(t)$ is the scale factor, dot stands for the time derivative here. The term $\psi=\dot{\phi}$ and $A_{,\phi}\equiv\frac{dA}{d\phi}$ (true for any function $A(\phi)$). By taking the time derivative of Eq. (2.3), we get
$12H\dot{H}\left(U_{0}-2\xi_{,\phi}H\psi\right)-12H^{2}\left(\xi_{,\phi\phi}H{\psi}^{2}+\xi_{,\phi}\left[H\dot{\psi}+\dot{H}\psi\right]\right)-\psi\dot{\psi}-V_{,\phi}\psi=0$ (2.6)
Expressing $V_{,\phi}$ from (2.6):
$V_{,\phi}=12H\dot{H}\left(\frac{U_{0}}{\psi}-2\xi_{,\phi}H\right)-12H^{2}\left(\xi_{,\phi\phi}H\psi+\xi_{,\phi}H\frac{\dot{\psi}}{\psi}+\xi_{,\phi}\dot{H}\right)-\dot{\psi}\,,$ (2.7)
and substituting it to (2.5), we obtain that Eq. (2.4) is a result of Eqs. (2.3) and (2.5).
When $\xi_{,\phi}\neq 0$, then equations (2.4) and (2.5) do not form a dynamic system. Insted it is proficient to use of Eqs. (2.4) and (2.5) t form a dynamical system [74]:
$\begin{split}\dot{\phi}=&\psi,\\ ~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\dot{\psi}=&\frac{1}{B-2\xi_{,\phi}H\psi}\left\{3\left[3-4\,\xi_{,\phi\phi}H^{2}\right]\xi_{,\phi}H^{2}\psi^{2}+\left[3B+2\,\xi_{,\phi}V_{,\phi}-6U_{0}\right]H\psi-\frac{V^{2}}{U_{0}}X\right\},\\ \dot{H}=&\frac{1}{4\left(B-2\,\xi_{,\phi}H\psi\right)}\left\{\left(4\,\xi_{,\phi\phi}H^{2}-1\right)\psi^{2}-16\xi_{,\phi}H^{3}\psi-4\frac{V^{2}}{U_{0}^{2}}\,\xi_{,\phi}H^{2}X\right\},\end{split}$ (2.8)
where
$B=12\xi_{,\phi}^{2}H^{4}+U_{0},\qquad X=\frac{U_{0}^{2}}{V^{2}}\left(12\xi_{,\phi}H^{4}+V_{,\phi}\right).$ (2.9)
Substituting $\dot{\psi}$ defined by Eq. (2.5) into Eq. (2.6), we get the third equation of system (2.8). So, this equation is a consequence of Eqs. (2.3) and (2.5).
As usual for inflationary model construction, the e-folding number $N=\ln(a/a_{e})$, where $a_{e}$ is a constant, is considered as a measure of time during inflation.
Using the relation ${d}/{dt}=H\,{d}/{dN}$ and introducing $\chi=\psi/H$, one can write system (2.8) as follows:
@@
where $Q\equiv H^{2}$.
@@
Equations (2.4) and (2.5) can be transformed to the following equations:
$2\frac{dQ}{dN}\left(U_{0}-2\xi_{,\phi}\chi Q\right)={}-Q\chi^{2}+4Q\left(\xi_{,\phi\phi}Q\chi^{2}+\xi_{,\phi}\left(\frac{\chi}{2}\frac{dQ}{dN}+Q\frac{d\chi}{dN}\right)-\xi_{,\phi}Q\chi\right),$ (2.14)
$\frac{\chi}{2}\,\frac{dQ}{dN}+Q\,\frac{d\chi}{dN}+3Q\,\chi={}-V_{,\phi}-6Q\,\xi_{,\phi}\left(\frac{dQ}{dN}+2Q\right).$ (2.15)
@@ III Reviewing PLP inflation
Beyond Poincare symmetry, supersymmetry is the last type of space-time symmetrisation permitted by the Coleman-Mandula no-go theorem[75]. According to [14], with this assumption, a superpotential may be expressed as:
$W=\frac{S(\phi_{1}-\phi_{2})}{2m}~{},$ (3.16)
where, $S,\phi_{1},\phi_{2}$ are the chiral superfields and $m$ is the associated sub-Planckian energy scale. From this idea, one can get the potential of the form:
@@
Where $M_{P}$ signifies the Planck mass and the $\phi$ is the inflation field. $V_{0}$ is the dimensionless parameter that needs to be fixed by matching the amplitude of the scalar power spectrum at the pivot scale to $2.098\times 10^{-9}$. Two potential parameters $n$ and $q$ can vary from $1$ to $4$, this leads to the $16$ different cases. Eq. 3.17 is the inflationary potential, we are interested in our study. It is obvious to see that, if $m$ is super-Planckian, then this leads to just a large field $\phi^{m}$ model class. Thus, keeping $m\leq 1$ in the Planck unit is necessary, for simplicity in our calculation, we take $m=1$. However, the analyses we present in this manuscript can be easily extended to other smaller values of $m$. The standard case has been studied in [14] and it needed a phase of thermal inflation to make it viable from the observational point of view. Studying such models in modified gravity can alleviate the necessity of thermal inflation.
@@ IV The slow-roll parameters without any approximation
Using system (2.10), we obtain that the parameter $\epsilon_{1}(N)$ satisfies the following equation:
@@
Dividing Eq. (2.4) on Eq. (2.3), we get another form of $\epsilon_{1}$:
$\varepsilon_{1}=\frac{3}{\psi^{2}+2V}\left[\psi^{2}-4H^{2}\left(\xi_{,\phi\phi}\psi^{2}+\xi_{,\phi}\dot{\psi}-H\xi_{,\phi}\psi\right)\right].$ (4.22)
Note that we do not use any approximation, so, these expressions for $\epsilon_{1}$ are exact.
It is useful, to rewrite evolution equations in terms of the slow-roll parameters. Equations (2.3) and (2.4) are equivalent to
$12U_{0}H^{2}\left(1-\delta_{1}\right)=\psi^{2}+2V=H^{2}\chi^{2}+2V\,,$ (4.23)
$4U_{0}\dot{H}\left(1-\delta_{1}\right)={}-\psi^{2}+2U_{0}H^{2}\delta_{1}\left(\delta_{2}+\varepsilon_{1}-1\right).$ (4.24)
From Eq. (4.24), it follows
$\psi^{2}=2U_{0}H^{2}\left[2\varepsilon_{1}-\delta_{1}+\delta_{1}\left(\delta_{2}-\varepsilon_{1}\right)\right]$ (4.25)
and
$\chi^{2}=2U_{0}\left[2\varepsilon_{1}-\delta_{1}+\delta_{1}\left(\delta_{2}-\varepsilon_{1}\right)\right]\,.$ (4.26)
@@
From Eq. (4.26), it follows that
$r=\left|\frac{U_{0}}{\xi_{,\phi}^{2}Q^{2}}\delta_{1}^{2}-8\delta_{1}\delta_{2}+8\delta_{1}\varepsilon_{1}\right|\,.$
After this, we can find
@@
Substituting expression (4.26) into Eq. (4.23), we get
$V=U_{0}H^{2}\left[6-2\varepsilon_{1}-5\delta_{1}-\delta_{1}\left(\delta_{2}-\varepsilon_{1}\right)\right]\,.$ (4.30)
Exact expression for $\frac{dQ}{dN}$ is
$\frac{dQ}{dN}=\frac{Q}{2\,U_{0}\left(1-\delta_{1}\right)}\left({}-\chi^{2}+2\,U_{0}\delta_{1}\left(\delta_{2}+\varepsilon_{1}-1\right)\right).$ (4.31)
@@ V New slow-roll approximations
To examine the stability of de Sitter solutions in model 2.1, an effective potential approch has been proposed in  [74] (see also, [45, 77]):
@@ V.1 New approximation I
Multiplying 4.23 to $Q$ and substituting $\psi$ in terms of the slow-roll parameter $\delta_{1}$, we obtain:
$12\,{U_{0}}\,\left(1-{\delta_{1}}\right)Q^{2}-2VQ-\frac{\delta_{1}^{2}U_{0}^{2}}{4\,\xi^{2}_{,\phi}}=0\,.$ (5.33)
Equation 5.33 always has a positive root $Q$ if $\delta_{1}<1$:
$Q=\frac{V}{12\,U_{0}\,\left(1-\delta_{1}\right)}+\frac{\sqrt{{V}^{2}\xi_{,\phi}^{2}+3\,U_{0}^{3}\delta_{1}^{2}\left(1-\delta_{1}\right)}}{12\,U_{0}\,\left(1-\delta_{1}\right)|\xi_{,\phi}|}\,.$ (5.34)
Let us assume that $\delta_{1}\ll 1$ and expand the obtained expression to series with respect to the slow-roll parameter $\delta_{1}$:
$Q\approx{\frac{V}{6\,U_{0}}}+{\frac{V}{6\,U_{0}}}\delta_{1}+{\cal{O}}(\delta^{2}_{1})\,.$ (5.35)
We construct slow-roll approximations with
@@
To distinguish the standard slow-roll approximation, we do not neglect $\delta_{1}$, so we should obtain $\delta_{1}(\phi)$ to get $Q(\phi)$. To do it we use the following approximation of Eq. (2.4):
$\dot{H}\simeq{}-\frac{\psi^{2}+4H^{3}\xi_{,\phi}\psi}{4U_{0}\left(1-\delta_{1}\right)}.$ (5.37)
Using Eq. (4.19) and neglecting terms proportional to $\delta_{1}^{3}$, we rewrite (5.37) as follows
$\dot{H}\simeq{}-\frac{\delta_{1}}{4(1-\delta_{1})}\left(\frac{U_{0}\delta_{1}}{4\xi_{,\phi}^{2}Q}+2Q\right)\simeq{}-\frac{Q\delta_{1}}{2}-\frac{U_{0}\delta_{1}^{2}}{16\xi_{,\phi}^{2}Q}-\frac{Q\delta_{1}^{2}}{2}\,.$ (5.38)
We neglect terms proportional to $\ddot{\phi}$ and ${\dot{\phi}}^{2}$ in Eq. (2.5) and use Eq. (4.19) to get the following approximate equation:
$\frac{3U_{0}\delta_{1}}{2\xi_{,\phi}}={}-V_{,\phi}-12Q\xi_{,\phi}\left(\dot{H}+Q\right).$ (5.39)
Substituting $Q$ and $\dot{H}$ from Eqs. (5.36) and (5.38) into Eq. 5.39 and neglecting terms, proportional to $\delta_{1}^{n}$, where $n\geqslant 2$, we get
@@
The knowledge of $\delta_{1}(\phi)$ allows us to obtain $Q(\phi)$ and $\chi(\phi)$. We obtain from Eq. (5.36) that
$Q\simeq\frac{V}{6U_{0}}\left[1-\frac{2V^{2}\xi_{,\phi}{V_{eff}}_{,\phi}}{V^{2}\xi_{,\phi}^{2}+3U_{0}^{3}}\right]=\frac{V\left(9U_{0}^{3}-6U_{0}^{2}\xi_{,\phi}\,V_{,\phi}+\xi_{,\phi}^{2}V^{2}\right)}{18U_{0}\left(3U_{0}^{3}+\xi_{,\phi}^{2}V^{2}\right)}.$ (5.41)
Therefore, Eq. (4.19) gives
$\chi=\frac{U_{0}\delta_{1}}{2\xi_{,\phi}Q}\simeq{}-\frac{6U_{0}^{2}V{V_{eff}}_{,\phi}}{V^{2}\xi_{,\phi}^{2}+3U_{0}^{3}-2V^{2}\xi_{,\phi}{V_{eff}}_{,\phi}}={}-\frac{6U_{0}^{2}\left(3U_{0}^{2}V_{,\phi}+\xi_{,\phi}V^{2}\right)}{V\left(9U_{0}^{3}-6U_{0}^{2}\xi_{,\phi}V_{,\phi}+\xi_{,\phi}^{2}V^{2}\right)}\,,$ (5.42)
hence,
Let us get the slow-roll parameters as functions of $\phi$. Using Eq. (5.41), we get
$\varepsilon_{1}(\phi)={}-\frac{1}{2}\frac{d\phi}{dN}\frac{d\ln(Q)}{d\phi}=\frac{3U_{0}^{2}\left(3U_{0}^{2}V_{,\phi}+\xi_{,\phi}V^{2}\right)}{V\left(9U_{0}^{3}-6U_{0}^{2}\xi_{,\phi}V_{,\phi}+\xi_{,\phi}^{2}V^{2}\right)}\,\frac{d\ln(Q)}{d\phi}\$ (5.44)

### v3

@@ Abstract
The results of Cosmic Microwave Background (CMB) observations reported by Planck satellite suggest that a plateau characteristic in the flat potential of the inflaton field is favored to drive the Universe’s early acceleration. On that note, Power Law Plateau (PLP) potential has gained a lot of attention. In this study, we demonstrate that implementing this inflationary model in an Einstein-Gauss-Bonnet(EGB) gravity background makes the model compatible with observations while avoiding late-stage thermal inflation, which is otherwise unavoidable in a standard scenario. More importantly, we have taken the PLP model as a case study to check the new slow-roll approximations method proposed in [1]. Then we take one step further to check the consequence of generalized reheating in this scenario. Thus opening a new promising window to study inflationary dynamics in EGB gravity.
@@ I Introduction
The inflationary epoch, along with the standard hot Big Bang model, could resolve the understanding of the birth and evolution of our Universe to its current state [2], [3]. The first idea of inflation was proposed by Guth [4], and since then, the inflationary epoch has gained tremendous attention, becoming an interdisciplinary study requiring both the fields of high energy physics and cosmology[5, 6, 7, 8, 9, 10, 11]. CMB observations by the Planck and Wilkinson Microwave Anisotropy Probe(WMAP) missions have ruled out or strongly disfavored most of the textbook models of inflation in the standard scenario[12, 13, 14]. Based on the current observation, it becomes evident that a plateau-like feature in the potential can satisfy the constraints imposed by CMB in the best possible manner. In a key study, Dimopoulos and Owen [15] presented the Power Law Plateau inflation (PLP) with plateau properties. However, in this theoretically justified model, one needs to incorporate a phase of thermal inflation to allow it to withstand the observational constraints while remaining in the conventional cold inflationary premise.
@@
It was shown for the first time in [16] that if one wants to develop the inflationary models under the UV complete string theory, PLP in the alternative theory of Randall-Sundrum brane-world[17, 18]can be compatible with the observation, and it can circumvent the tighter constraints from string theory, such as swampland conjecture[19]. However, on the other hand, RS models suffer serious existential crises because of the continuous limitation on the scale provided by the braneworld theory[20]. It has been shown that in modified gravity, the restriction on the inflationary models can be less stringent[21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39]. In other words, models that are disfavoured by observations in the standard cold inflationary model can satisfy the observational constraints in the modified gravity. The addition of the Gauss-Bonnet component to the Einstein-Hilbert action does not affect the equations of motion, as it is a total derivative. When EGB coupling $\xi(\phi)$ is combined with a function of the scalar field $\phi$, it becomes dynamically important. In string theory, the Gauss-Bonnet term serves as a quantum correction to the Einstein-Hilbert action. Numerous inflationary models have been studied in this context[40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 57, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83]; the most widely studied EGB coupling is the inverse function of the scalar field[41, 42, 43, 44, 46, 54, 56, 59, 60]. However, in this manuscript, we will study a different form of coupling that involves a $\tanh$" term[84]. In [1], the authors have presented a new kind of slow-roll approximations, which are found to be more accurate in comparison to the previous slow-roll approximations presented in[50]. As it has been already shown in [1] that new slow approximations are more accurate, we aim to do a case study of the new slow-roll approximations in the EGB scenario using the the PLP potential as a prototype. This serves the primary objective of this exercise. More importantly, we extend the study to accommodate the generalized reheating and review its effect in the effective EGB background. The standard cold inflationary scenario requires an epoch of reheating after the end of inflation to enter the radiation-dominated epoch and start the Big Bang Nucleosynthesis (BBN). Studying the reheating phase following the end of inflation is necessary for the conceptual development of inflationary model construction. We have studied reheating in this work and established the limits on the number of e-folding and reheating temperatures that are permitted by CMB data.
@@
The paper is organized as follows: In section II, we will go over some of the fundamental critical features in the EGB framework. A theoretically well-motivated Power Law Plateau (PLP) potential for this study has been reviewed in section III. In section IV, the calculation of slow-roll parameters without any approximations is reviewed. In section V, the new slow-roll approximations (SRA) as reported in the work [1] are reviewed. Finally, the analysis of the inflationary dynamics using the new SRA and the effect of generalized reheating has been reported in the sections VI and VII. Finally, we conclude in section VIII with our conclusions and future direction of work.
@@ II Review of Einstein-Gauss Bonnet Gravity
The evolution of equations and dynamical systems in Einstein-Gauss-Bonnet gravity can be written for a spatially flat FLRW metric [40, 50]as:
@@
where $H=\dot{a}/a$ denotes the Hubble parameter and $a(t)$ is the scale factor, the dot stands for the time derivative here. The term $\psi=\dot{\phi}$ and $A_{,\phi}\equiv\frac{dA}{d\phi}$ (true for any function $A(\phi)$).
Using the relation d/dt=Hd/dN{d}/{dt}=H\,{d}/{dN} and introducing $\chi=\psi/H$, one can write system. (For detailed calculation see [1]):
@@
where $Q\equiv H^{2}$, $B=12\xi_{,\phi}^{2}H^{4}+U_{0}$ and $X=\frac{U_{0}^{2}}{V^{2}}\left(12\xi_{,\phi}H^{4}+V_{,\phi}\right)$
@@ III Reviewing PLP inflation
As the most advanced extension of space-time symmetries permitted by the Coleman–Mandula no-go theorem [85], supersymmetry goes beyond Poincaré symmetry. Within this setting, and as shown in [15], one can write the superpotential as;
$W=\frac{S^{2}(\Phi^{2}_{1}-\Phi^{2}_{2})}{2m}~,$ (3.10)
Here, $S,\Phi_{1},\Phi_{2}$ are the chiral superfields, while $m$ represents the corresponding sub-Planckian energy scale. From [15], the F-term scalar potential takes the form:
$V_{F}=\frac{|S|^{2}}{m^{2}}\Big{[}|\Phi^{2}_{1}-\Phi^{2}_{2}|^{2}+|S|^{2}\left(|\Phi_{1}|^{2}+|\Phi_{2}|^{2}\right)\Big{]}~.$ (3.11)
The potential 3.11 gets minimised for $\Phi_{1}=\Phi_{2}$ . After imposing an appropriate R-symmetry and performing the field rotation in the configuration space, we can introduce the canonically normalised real scalar field $\phi$ such as $|\Phi_{1}|=|\Phi_{2}|\equiv\frac{1}{2}\phi$. This leads to the following;
$V_{F}=\frac{|S|^{4}\phi^{2}}{2m^{2}}.$ (3.12)
After considering the D-term contribution to the scalar potential [86], we are left with;
$V_{D}=\frac{1}{2}\left(|S|^{2}-\sqrt{2}M^{2}\right)^{2}\;,$ (3.13)
where $M$ signifies the scale of grad unified theory (GUT), the total potential can be denoted as:
$V=\frac{|S|^{4}\phi^{2}}{2m^{2}}+\frac{1}{2}\left(|S|^{2}-\sqrt{2}M^{2}\right)^{2}~.$ (3.14)
After imposing the condition that the potential is minimized with respect to the S direction
$\frac{\partial V}{\partial|S|}=0\;\Rightarrow\;\langle|S|^{2}\rangle=\frac{\sqrt{2}M^{2}}{1+\phi^{2}/m^{2}}~.$ (3.15)
Substituting Eq. 3.15 into Eq. 3.14 we get:
$V=\frac{M^{4}\phi^{2}}{m^{2}+\phi^{2}}~,$ (3.16)
and writing $M^{4}=V_{0}M^{4}_{P}$, we get the following form of the potential
$V{(\phi)}=V_{0}M_{P}^{4}\frac{\phi^{2}}{m^{2}+\phi^{2}}~.$ (3.17)
Following [15] we obtained a more generalized form of the potential:
@@
where $M_{P}$ signifies the Planck mass and the $\phi$ is the inflation field. $V_{0}$ is the dimensionless parameter that needs to be fixed by matching the amplitude of the scalar power spectrum at the pivot scale to $2.0989\times 10^{-9}$. Two potential parameters $n$ and $q$ can vary from $1$ to $4$, leading to $16$ different cases. Eq. (3.18) is the inflationary potential, which we are interested in for our study. It is obvious to see that, if $m$ is super-Planckian, then this leads to just a large field $\phi^{m}$ model class. Thus, keeping $m\leq 1$ in the Planck unit is necessary, for simplicity in our calculation, we take $m=1$. However, the analyses we present in this manuscript can easily be extended to other smaller values of $m$. The standard case has been studied in [15], and it required a phase of thermal inflation to make it viable from the observational point of view. Studying such models in modified gravity can alleviate the necessity of thermal inflation.
@@ IV The slow-roll parameters without any approximation
Using system (2.6), we obtain that the parameter $\varepsilon_{1}(N)$ satisfies the following equation:
@@
After this, we can find the amplitude of the scalar power spectrum as:
@@ V New slow-roll approximations
To examine the stability of de Sitter solutions in model (2.1), an effective potential approach has been proposed in  [88] (see also, [50, 89]):
@@ V.1 New approximation I
We construct slow-roll approximations with. (For detailed calculation see [1])
@@
Following  [1] we write the slow roll parameters as:
@@
$\varepsilon_{1}(\phi)={}-\frac{1}{2}\frac{d\phi}{dN}\frac{d\ln(Q)}{d\phi}=\frac{3U_{0}^{2}\left(3U_{0}^{2}V_{,\phi}+\xi_{,\phi}V^{2}\right)}{V\left(9U_{0}^{3}-6U_{0}^{2}\xi_{,\phi}V_{,\phi}+\xi_{,\phi}^{2}V^{2}\right)}\,\frac{d\ln(Q)}{d\phi}\\$ (5.29)
@@
The relation between the number of e-folds andthe field can be written as [1]: